OpenStudy (anonymous):

\[\int\limits \frac{ secxtanx }{ 4+\sec^2x }\]

OpenStudy (anonymous):

im pretty much lost with the trig functions

OpenStudy (anonymous):

i know i have to use inverse functions but secxtanx, is the inverse just secx?

OpenStudy (shubhamsrg):

a simple substitution of secx = t will help your case.

OpenStudy (anonymous):

could you show me?

OpenStudy (shubhamsrg):

secx = t secx tanx dx = dt rest is easier

OpenStudy (anonymous):

so is it u/4+sec^2x ?

OpenStudy (shubhamsrg):

nop o.O it becomes dt/(4 + t^2)

OpenStudy (anonymous):

\[\frac{ secxtanx }{ 4 + (secxtanx)^2 }= 1/4 +\sec x tanx\]

OpenStudy (anonymous):

is that my answer or am i still way off?